OpenStudy (anonymous):

the perimeter of a rectangular is 162 m . the length is 9m more then twice the width find the dimensions. what is the length?

OpenStudy (blacksteel):

I posted this in your other post with the same question, but in case you're looking here: The perimeter is twice the length plus twice the width. If the length is L and the width is W, the perimeter P is P = 2W + 2L. we also know that P = 162, so 2W + 2L = 162. We know that the length is twice the width + 9, so L = 2W + 9. Then we can substitute this answer into the perimeter equation to get 2W + 2(2W + 9) = 162. Simplifying this gives us 6W + 18 = 162, or 6W = 144. Dividing both sides by 6 gives us W = 24. Since we now know W, we can substitute this back into our second equation, L = 2W + 9. This gives us L = 2(24) + 9, or L = 57. (Saifoo's technique was correct, but the problem reads that the length is twice the width + 9, not the width + 9.)

OpenStudy (anonymous):

i give up on this questions

OpenStudy (anonymous):

P=(L+W)2 L=9+2W 162=((9+2W)+W))2 162=(9+3W)2 162=18+6W 162-18=6W 144 /6=W 24 = Width length = 9 +2(24) =9+48 length = 57

OpenStudy (anonymous):

From the diagram: \[6w + 18 = 162 \rightarrow w = \frac{144}{6} = 24\] so \[l = 2w + 9 = 57\]